# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v1

@@ 4.3. When Wi-Fi Sensing Meets Large Multimodal Models
  3. (3)
    How should domain-specific knowledge be incorporated into the model? Wi-Fi sensing encompasses a wealth of domain-specific knowledge, including signal propagation principles, communication protocols, e.g. IEEE 802.11 bf, and established research findings. Should this information be used to fine-tune models before fine-tuning on Wi-Fi signals, or would integrating Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) as an external database allow the model to access Wi-Fi knowledge on demand? How can this domain knowledge improve interpretability, efficiency, and overall performance?
  4. (4)
@@
4.4. System Deployment and Service Models
@@
  1. (1)
    Large multimodal models often require high computational power and storage. Compression techniques like pruning (Han et al., 2015), quantization (Jacob et al., 2018), or knowledge distillation (Hinton et al., 2015) can help balance performance and efficiency. How can these methods be tailored for different hardware (e.g., edge devices or mobile phones) to ensure speed without sacrificing too much performance?
  2. (2)
    How to manage network bandwidth, data traffic, and latency during operation? Network constraints, particularly with real-time data transmission, pose challenges. How can continuous data flow between edge devices and servers be optimized? Strategies like edge computing (Shi et al., 2016), which processes data locally, could reduce latency and bandwidth usage but introduce complexity in resource management and synchronization.
  3. (3)
@@
4.5. Security and Privacy
The integration of Wi-Fi sensing in various applications introduces several security and privacy concerns. These issues need to be addressed to ensure the safe and ethical use of Wi-Fi sensing technologies in real-world scenarios. Some of the key challenges include:
  1. (1)
    How can sensitive user data be handled effectively? Wi-Fi sensing data may inadvertently capture sensitive information, such as user identity or pose (Wang et al., 2019b, f). How can we ensure that such details are not exposed in applications like crowd counting, where they are unnecessary? What strategies can be employed to anonymize or remove sensitive data without compromising the performance of task-specific models? Can techniques such as differential privacy, data anonymization, or task-specific filtering be used to ensure that only relevant information is retained while protecting user privacy?
  2. (2)
    How can sensing models be made resilient to attacks? Sensing models may be vulnerable to adversarial attacks that disrupt or manipulate the system output (Liu et al., 2022). How can we protect Wi-Fi sensing systems from attackers who try to interfere with the communication channel and alter the sensing results? What approaches, such as robust model training, error correction, and secure data transmission protocols, can be adopted to mitigate such risks and ensure the reliability of applications?
  3. (3)
    What privacy concerns arise from Wi-Fi penetration ability? Wi-Fi sensing has the potential to see through walls (Adib and Katabi, 2013) and capture the behavior of individuals in adjacent rooms or neighboring environments. How can we address privacy leakage concerns, particularly in residential or private spaces? What privacy-preserving protocols need to be designed to limit the scope of sensing to the intended areas? Should regulations or hardware modifications be implemented to restrict the range of Wi-Fi signals and ensure that sensitive information is not accessed without consent?
In conclusion, as Wi-Fi sensing technologies advance, it is imperative to integrate robust security and privacy measures that protect users and ensure the ethical use of the technology. Developing secure models, using privacy-preserving techniques, and complying with relevant regulations will be key to the success and wide adoption of Wi-Fi sensing applications.
5. Conclusion
This survey encompasses a significantly broader range of Wi-Fi generalization studies—over 200 papers in total. We have organized these studies along the entire Wi-Fi sensing pipeline: from device deployment and signal processing to feature learning and model deployment. This structured approach enables researchers to quickly locate the techniques that best address their specific challenges. Furthermore, we provide a detailed overview of open-source datasets, helping researchers efficiently identify the data needed to validate their algorithms. Finally, we offer a comprehensive analysis of future directions, including trends in multimodal approaches and large language models, to inspire new research ideas. In summary, this work serves as a valuable resource for researchers seeking convenient access to methodologies, datasets, and innovative research strategies.
Contributions
Fei Wang initiated and led the project, reviewed and organized all the papers on Wi-Fi generalizability and datasets, and was responsible for writing the manuscript. Tingting Zhang reviewed and organized approximately one-fifth of the generalizability and dataset papers, participating in discussions on the manuscript. Bintong Zhao, Libao Xing, and Tiantian Wang each reviewed and categorized approximately one-fifth of the Wi-Fi generalizability papers.

### v2

@@ IV-C When Wi-Fi Sensing Meets Large Multimodal Models
  3. 3.
    How should domain-specific knowledge be incorporated into the model? Wi-Fi sensing encompasses a wealth of domain-specific knowledge, including signal propagation principles, communication protocols, e.g. IEEE 802.11 bf, and established research findings. Should this information be used to fine-tune models before fine-tuning on Wi-Fi signals, or would integrating Retrieval-Augmented Generation (RAG) [318] as an external database allow the model to access Wi-Fi knowledge on demand? How can this domain knowledge improve interpretability, efficiency, and overall performance?
  4. 4.
@@
IV-D System Deployment and Service Models
@@
  1. 1.
    Large multimodal models often require high computational power and storage. Compression techniques like pruning [319], quantization [320], or knowledge distillation [321] can help balance performance and efficiency. How can these methods be tailored for different hardware (e.g., edge devices or mobile phones) to ensure speed without sacrificing too much performance?
  2. 2.
    How to manage network bandwidth, data traffic, and latency during operation? Network constraints, particularly with real-time data transmission, pose challenges. How can continuous data flow between edge devices and servers be optimized? Strategies like edge computing [322], which processes data locally, could reduce latency and bandwidth usage but introduce complexity in resource management and synchronization.
  3. 3.
@@
V Conclusion
In this survey, we systematically reviewed and categorized Wi-Fi sensing generalization studies published between 2014 and 2025, following the end-to-end pipeline of Wi-Fi sensing systems—spanning device deployment, model construction, and model deployment. We summarized key techniques proposed in each stage to address the challenges of domain variability. In addition, we provided a comprehensive overview of existing publicly available datasets, highlighting their domain diversity and applicability to generalization research.
By integrating methodological insights with data resources, this survey serves as a practical handbook for researchers aiming to advance the generalizability of Wi-Fi sensing systems. Beyond the current landscape, we also discussed future research directions, including synthetic dataset generation, pretraining large-scale perception models, integration with multimodal foundation models, and deployment-aware continual learning.
[…]